# On the Optimization Dynamics of RLVR: Gradient Gap and Step Size Thresholds

### v2

@@ 5 Numerical Experiments – REINFORCE on Contextual Bandits
We consider a contextual variant of the synthetic bandit experiment of (arnal2025asymmetric, , Section 5.1). contextual bandit with contexts $x\in[0,1]^{d}$ for $d=10$. For a set of N:=100N:\,=100 arms, we generate linear scores for each context $x$, ${\bf s}(x)=\boldsymbol{\beta}^{\top}x\in\mathds{R}^{N}$ for a matrix $\boldsymbol{\beta}\in\mathds{R}^{d\times N}$, with standard normal entries. $r_{y}(x):\,=\operatorname*{arg\,max}_{y\in[N]}{\bf s}(x)$. We use linear logits initialized as $\ell_{0}(x):\,=\theta_{0}^{\top}x\in\mathds{R}^{N}$, for $\theta_{0}\sim\mathcal{N}(0,0.01^{2}\cdot\text{Id}_{d\times d})$. The policy $\uppi_{\theta_{k}}$ was then initialized as a softmax over $\ell_{0}$ and the parameters $\theta_{k}$ were updated according to the REINFORCE exact gradient update at a training context $x_{k}$ with stepsize $\eta:\,=0.1$:
$\theta_{k+1}=\theta_{k}+\eta\,\,\mathds{E}_{y\sim\uppi_{\theta_{k}}(\cdot\mid x_{k})}[(r_{y}(x_{k})-J_{x_{k}}(\uppi_{\theta_{k}}))\cdot\nabla_{\theta}\,\log\uppi_{\theta_{k}}(y\mid x_{k})]\,.$
@@
The training context $x_{k}$ was selected at random among those (from an initial pool of $100$ contexts drawn uniformly from $0,1]^{d}$) with intermediate value function $J(x_{k})\in[0.2,0.8]$, following intuitions from curriculum learning for filtering out overly difficult or easy prompts (zhang2025speedrl, ).
@@
We construct three plots based on calculating the following for $500$ randomly evaluated contexts $x$: the value function $J_{x}(\uppi_{\theta_{k}})$, per-context cumulative gradient gap $\sum_{i=0}^{k}[\Delta\mu_{x}(i)]_{+}\cdot\eta$, and the relative per-context cumulative gradient gap $\sum_{i=0}^{k}([\Delta\mu_{x}(i)]_{+}-[\Delta\mu_{x_{i}}(i)]_{+})\cdot\eta$ which measures the discrepancy of the gradient gaps at the training contexts $x_{k}$.
@@
From the first subplot, we observe a distinctive logistic relationship between the cumulative gradient gap and the value function, reminiscent of our theory (Corollary˜1).
From the second subplot, we see there are two regimes for each context’s cumulative gradient gap curve: either fast exponential convergence (Corollary˜1) or lack of improvement (Theorem˜2).
In the third plot, we interestingly see that those contexts with close to 0 relative cumulative gradient gap (i.e., close to that of training contexts) experience faster convergence.
@@ 6 Discussion and Future Directions
Our analysis is restricted to the single-prompt setting, which enabled sharp characterizations of Gradient Gap alignment and step size scaling. In practice, however, training involves a diverse batch of prompts. In this regime, both the alignment signal $\Delta\mu_{q}(k)$ and the optimal step size $\eta_{k}$ can vary substantially across prompts. A single update direction $\boldsymbol{w}_{k}$ may align well with some prompts but poorly with others, and a step size that is safe for one subset may be overly aggressive for another, leading to overshooting and limited overall gains.
These observations suggest several directions for future work: developing prompt-adaptive updates that adjust direction or scale based on batch heterogeneity, analyzing the statistical dynamics of RLVR under diverse prompt distributions, and extending the framework to sequential or curriculum-based training (bengio2009curriculum, ; chen2025selfevolving, ; zhang2025speedrl, ). Such extensions are essential for a full theory of RLVR in realistic multi-prompt settings.
@@ Appendix A Additional Related Work
A growing body of work has begun to examine the theoretical foundations of preference-based RLHF and verifiable-reward RL. Early studies analyzed preference-driven RL with trajectory-level feedback, establishing convergence guarantees under pairwise or $K$-wise comparisons [15, 8, 30]. More recent results offered complexity characterizations: Want et al. [22] and Du et al. [10] compared RLHF with standard RL, identifying conditions for sample-efficient preference optimization. In parallel, techniques from function approximation and offline RL have been adapted to the fine-tuning setting [6, 21, 5]. Yet the optimization behavior of Reinforcement Learning with Verifiable Rewards (RLVR)—where supervision is provided by deterministic outcomes rather than preferences—remains largely unexplored.
Building on these foundations, researchers have investigated how sparse, outcome-based rewards shape learning dynamics. For example, one study shows that off-policy updates can benefit from emphasizing positive (successful) outcomes more strongly than negative ones [2], while another finds that purely outcome-driven signals can collapse solution diversity in the absence of exploration incentives [18]. These insights illustrate how verifiable binary rewards influence both convergence and the diversity of reasoning strategies, offering early theoretical guidance for RLVR.
Complementing this theoretical line, sequence-level optimization methods have advanced the algorithmic toolkit for RL-based fine-tuning. GSPO, for instance, defines importance ratios over whole-answer likelihoods with sequence-level clipping and updates, improving stability and efficiency compared to token-level methods [29]. To control variance from variable output lengths, specialized loss aggregation schemes such as $\Delta L$ normalization have been introduced, yielding an unbiased, minimal-variance estimator of policy loss across different generation lengths [12].
@@ C.1.1 Proof of Theorem 1
We begin with the following Lemma˜1, which is the core tool for our analysis. The proof is provided in Section˜E.1.
@@
Suppose that ˜1 holds and the step size
@@
Iterating inequality (28) over $T$ timesteps gives us a bound on the cumulative effect:
@@
First, we show that return $J_{q}(K)$ can be bounded away from the optimum under the condition from Theorem˜1(a). Rearranging the cumulative inequality gives an upper bound:
@@
By applying the bounds on the cumulative alignment gap and step sizes from Theorem˜1(a), the inequality simplifies to
@@
This result shows that the performance $J_{q}(K)$ hits a ceiling and is strictly bounded away from the optimum value of $1$, proving the stagnation described in Theorem˜1(a).
@@
This result shows that as the cumulative gap alignment grows, the exponential term shrinks, pushing the performance $J_{q}(K)$ towards $1$. This confirms the convergence guarantee in Theorem˜1(b), as described in inequality (15b).
C.1.2 Proof of Corollary˜1
@@
Next, rearranging and telescoping in the same manner as the proof of Theorem˜1, we obtain
@@
which completes the proof of Corollary˜1.
@@ C.2 Token-Level Analysis (Proof of Theorem 3)
We begin by providing the formal statement of inequality (22) from the main body. This result, presented below as Lemma˜2, serves as the token-level counterpart to the trajectory-level analysis in Lemma˜1 and establishes a precise lower bound on the one-step improvement of the log-odds ratio.
@@
With this lemma, we can proceed to the main proof of Theorem˜3. Our strategy is to synthesize the bounds from Lemma˜1 (the trajectory-level analysis) and Lemma˜2 (the token-level analysis). Our central aim is to demonstrate that for a sufficiently small step size $\eta_{k}$, the log-odds of the objective function improves at each step according to the following inequality:
@@
In this scenario, we apply the token-level bound from Lemma˜2. The lemma guarantees that inequality (32) holds, provided that the step size $\eta_{k}$ meets conditions
@@
For the alternative case, we fall back on the trajectory-level bound from Lemma˜1. By setting the Lipschitz and gradient norm parameters to $L_{\mathrm{o}}=L_{\mathrm{p}}T_{\infty}$ and $G_{\mathrm{o}}=G_{\mathrm{p}}T_{\infty}$ respectively, the bound (28) ensures that inequality (32) holds if
@@
With this per-step improvement established, the remainder of the proof follows the same structure as the proof of Theorem˜1(b). We sum inequality (32) over all iterations from $k=0$ to K1K-1, which yields the final convergence result (18b) and completes the proof.
Appendix D Proofs of Lower Bounds
D.1 Trajectory-Level Analysis Lower Bound (Proof of Theorem 2)

### v3

@@ 6.1 REINFORCE on Contextual Bandits
We consider a contextual variant of the synthetic bandit experiment of Arnal et al. (2025, Section 5.1): a contextual bandit with contexts $x\in[0,1]^{d}$ for $d=10$. For a set of N:=100N:\,=100 arms, we generate linear scores for each context $x$, ${\bf s}(x)=\boldsymbol{\beta}^{\top}x\in\mathds{R}^{N}$ for a matrix $\boldsymbol{\beta}\in\mathds{R}^{d\times N}$, with standard normal entries, and one-hot encoded reward $r_{y}(x):\,=\mathds{1}\{y\in\operatorname*{arg\,max}_{y\in[N]}{\bf s}(x)\}$. We use linear logits initialized as $\ell_{0}(x):\,=\theta_{0}^{\top}x\in\mathds{R}^{N}$, for $\theta_{0}\sim\mathcal{N}(0,0.01^{2}\cdot\text{Id}_{d\times d})$. The policy $\uppi_{\theta_{k}}$ was then initialized as a softmax over $\ell_{0}$ and the parameters $\theta_{k}$ were updated according to the REINFORCE exact gradient update at a training context $x_{k}$ with stepsize $\eta:\,=0.1$:
$\theta_{k+1}$ $=\theta_{k}+\eta\,\,\mathds{E}_{y\sim\uppi_{\theta_{k}}(\cdot\mid x_{k})}\big[(r_{y}(x_{k})-J_{x_{k}}(\uppi_{\theta_{k}}))\cdot\nabla_{\theta}\,\log\uppi_{\theta_{k}}(y\mid x_{k})\big]\,.$
@@
The training context $x_{k}$ was selected at random among those (from an initial pool of $100$ contexts drawn uniformly from $0,1]^{d}$) with intermediate value function $J(x_{k})\in[0.2,0.8]$, following intuitions from curriculum learning for filtering out overly difficult or easy prompts (Zhang et al., 2025).
@@
We construct three plots based on calculating the following for $500$ randomly evaluated contexts $x$: the value function $J_{x}(\uppi_{\theta_{k}})$, per-context cumulative Gap Alignment $\sum_{i=0}^{k}[\Delta\mu_{x}(i)]_{+}\cdot\eta$, and the relative per-context cumulative Gap Alignment $\sum_{i=0}^{k}([\Delta\mu_{x}(i)]_{+}-[\Delta\mu_{x_{i}}(i)]_{+})\cdot\eta$ which measures the discrepancy of the Gap Alignments at the training contexts $x_{k}$.
@@
From the first subplot, we observe a distinctive logistic relationship between the Cumulative Gap Alignment and the value function, reminiscent of our theory (Corollary 3.1). From the second subplot, we see there are two regimes for each context’s Cumulative Gap Alignment curve: either fast exponential convergence (Corollary 3.1) or lack of improvement (Theorem 2). In the third plot, we interestingly see that those contexts with close to 0 relative Cumulative Gap Alignment (i.e., close to that of training contexts) experience faster convergence.
6.2 GRPO on Language Models
We validate our theory on three GRPO training runs on the base model Qwen2.5-Math-7B for three math reasoning datasets: (1) the GSM8K dataset (Cobbe et al., 2021), (2) the DeepScaleR dataset (Luo et al., 2025), and (3) the DAPO-17k dataset (Yu et al., 2025). For background, the GSM8K dataset consists of grade-school math word problems, while the more challenging DeepScaleR and DAPO-17k datasets consist of problems derived from past AIME and AMC competitions.
At each training step, we approximate the batch-average Alignment Gap $\mathds{E}_{q}[\Delta\mu_{q}]$ using the relation $\boldsymbol{g}_{\mathrm{GRPO}}\propto\sqrt{J_{q}(1-J_{q})}\cdot\big(\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}\big)$, as derived in Section 4. In Figure 2, we plot the Cumulative Gap Alignment vs. the value function, colored by normalized step count. For all three datasets, we see a similar relationship between Cumulative Gap Alignment and accuracy as in our theory and bandit experiment.
Figure 2: Cumulative Gap Alignment vs. Accuracy for our experiments.
@@ 7 Discussion and Future Directions
We sharply characterized the Gradient Gap alignment and step size scaling for policy gradient methods. In practice, the per-prompt alignment signal $\Delta\mu_{q}(k)$ and subsequent optimal step size $\eta_{k}$ can vary substantially across prompts $q$. A single update direction $\boldsymbol{w}_{k}$ may align well with some prompts but poorly with others, and a step size that is safe for one subset may be overly aggressive for another, leading to overshooting and limited overall gains.
These observations suggest several directions for future work: developing prompt-adaptive updates that adjust direction or scale based on batch heterogeneity, analyzing the statistical dynamics of RLVR under diverse prompt distributions, and extending the framework to sequential or curriculum-based training (Bengio et al., 2009; Chen et al., 2025b; Zhang et al., 2025).
@@ Appendix A Additional Related Work
A growing body of work has begun to examine the theoretical foundations of preference-based RLHF and verifiable-reward RL. Early studies analyzed preference-driven RL with trajectory-level feedback, establishing convergence guarantees under pairwise or $K$-wise comparisons (Pacchiano et al., 2021; Chen et al., 2022; Zhu et al., 2023). More recent results offered complexity characterizations: Wang et al. (2023) and Du et al. (2024) compared RLHF with standard RL, identifying conditions for sample-efficient preference optimization. In parallel, techniques from function approximation and offline RL have been adapted to the fine-tuning setting (Chen, 2025; Wang et al., 2024; Brantley et al., 2025). Yet the optimization behavior of Reinforcement Learning with Verifiable Rewards (RLVR)—where supervision is provided by deterministic outcomes rather than preferences—remains largely unexplored.
Building on these foundations, researchers have investigated how sparse, outcome-based rewards shape learning dynamics. For example, one study shows that off-policy updates can benefit from emphasizing positive (successful) outcomes more strongly than negative ones (Arnal et al., 2025), while another finds that purely outcome-driven signals can collapse solution diversity in the absence of exploration incentives (Song et al., 2025). These insights illustrate how verifiable binary rewards influence both convergence and the diversity of reasoning strategies, offering early theoretical guidance for RLVR.
Complementing this theoretical line, sequence-level optimization methods have advanced the algorithmic toolkit for RL-based fine-tuning. GSPO, for instance, defines importance ratios over whole-answer likelihoods with sequence-level clipping and updates, improving stability and efficiency compared to token-level methods (Zheng et al., 2025). To control variance from variable output lengths, specialized loss aggregation schemes such as $\Delta L$ normalization have been introduced, yielding an unbiased, minimal-variance estimator of policy loss across different generation lengths (He et al., 2025).
@@ C.1.1 Proof of Theorem 1
We begin with the following Lemma 1, which is the core tool for our analysis. The proof is provided in Subsection D.1.
@@
Suppose that Assumption 1 holds and the step size
@@
Iterating inequality Equation 27 over $K$ timesteps gives us a bound on the cumulative effect:
@@
First, we show that the return $J_{q}(K)$ can be bounded away from the optimum under the condition from Theorem 1(a). Rearranging the cumulative inequality gives an upper bound:
@@
By applying the bounds on the Cumulative Gap Alignment and step sizes from Theorem 1(a), the inequality simplifies to
@@
This result shows that the performance $J_{q}(K)$ hits a ceiling and is strictly bounded away from the optimum value of $1$, proving the stagnation described in Theorem 1(a).
@@
This result shows that as the Cumulative Gap Alignment grows, the exponential term shrinks, pushing the performance $J_{q}(K)$ towards $1$. This confirms the convergence guarantee in Theorem 1(b), as described in inequality (13b).
C.1.2 Proof of Corollary 3.1
@@
Next, rearranging and telescoping in the same manner as the proof of Theorem 1, we obtain
@@
which completes the proof of Corollary 3.1.
@@ C.2 Token-Level Analysis (Proof of Theorem 3)
We begin by providing the formal statement of inequality (19) from the main body. This result, presented below as Lemma 2, serves as the token-level counterpart to the trajectory-level analysis in Lemma 1 and establishes a precise lower bound on the one-step improvement of the log-odds ratio. The proof is found in Subsection D.2.
@@
With this lemma, we can proceed to the main proof of Theorem 3. Our strategy is to synthesize the bounds from Lemma 1 (the trajectory-level analysis) and Lemma 2 (the token-level analysis). Our central aim is to demonstrate that for a sufficiently small step size $\eta_{k}$, the log-odds of the objective function improves at each step according to the following inequality:
@@
In this scenario, we apply the token-level bound from Lemma 2. The lemma guarantees that inequality (31) holds, provided that the step size $\eta_{k}$ meets conditions
@@
For the alternative case, we fall back on the trajectory-level bound from Lemma 1. By setting the Lipschitz and gradient norm parameters to $L_{\mathrm{o}}=L_{\mathrm{p}}T_{\infty}$ and $G_{\mathrm{o}}=G_{\mathrm{p}}T_{\infty}$ respectively, the bound (27) ensures that inequality (31) holds if
@@
With this per-step improvement established, the remainder of the proof follows the same structure as the proof of Theorem 1(b). We sum inequality (31) over all iterations from $k=0$ to K1K-1, which yields the final convergence result (15b) and completes the proof.
Appendix D First-Order Analysis of the Log-Odds Change (Proofs of Auxilliary Lemmas for Convergence Guarantees)
Trajectory Level:
The analysis for entire generation trajectories is presented in Lemma 1, with the full proof in Subsection D.1.
Token Level:
The corresponding result for token-level consideration is established in Lemma 2, with its proof located in Subsection D.2.
D.1 Proof of Lemma 1
D.1.1 Overview
We first formulate Lemma 3 below, which establishes an equivalent expression for the difference of logarithms. The proof is deferred to Subsection D.1.2.
Lemma 3.
For policies $\uppi_{\theta_{\rm old}}$ and $\uppi_{\theta}$, it holds that
The difference of logarithms from Lemma 3 is central to our analysis of the optimization scheme. As shown in equation (34), we express this term as the difference between two log moment-generating functions (MGFs). These MGFs are derived from the conditional probabilities over the correct (positive) and incorrect (negative) solution spaces, $\mathcal{O}_{q}^{+}$ and $\mathcal{O}_{q}^{-}$, respectively.
In Lemma 4 below, we approximate the log MGFs based on the expectations of the random variables. See Subsection D.1.3 for the proof.
Lemma 4.
For any random variable $X$, we have:
We now combine Lemmas 3 and 4 to prove Lemma 1. In what follows, we introduce a shorthand $\boldsymbol{L}_{\theta}:\,=\log\uppi_{\theta}$. In our analysis recall that $\theta-\theta_{\rm old}=\eta_{k}\cdot\boldsymbol{w}_{k}$ with $\|\boldsymbol{w}_{k}\|_{2}\leq 1$. We denote $\eta:\,=\eta_{k}$ and $\boldsymbol{w}:\,=\boldsymbol{w}_{k}$ for short.
We apply the bound from equation (35) in Lemma 4 to the random variable $(\boldsymbol{L}_{\theta}-\boldsymbol{L}_{\theta_{\rm old}})(q,\vec{\boldsymbol{o}}\!\,)$. By evaluating this for the distributions corresponding to the positive space, $\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta_{\rm old}}\big(\cdot\bigm|q,\mathcal{O}_{q}^{+}\big)$, and the negative space, $\vec{\boldsymbol{o}}\!\,\sim\uppi_{\theta_{\rm old}}\big(\cdot\bigm|q,\mathcal{O}_{q}^{-}\big)$, we find that
where
To establish the final bound in Lemma 1, we need to prove two intermediate results:
  1. 1.
    $T_{1}=\Delta\mu_{q}(\uppi_{\theta_{\rm old}})\cdot\eta+\mathcal{O}(\eta^{2})$
  2. 2.
Analyzing term $T_{1}$:
For term $T_{1}$, Assumption 1 on the smoothness of function $\boldsymbol{L}_{\theta}$ implies that
Moreover, the definition (10) of gap $\Delta\mu_{q}$ leads to
From inequality (37), we also have
Finalizing the proof:
Combining our previous bounds (36), (38) and (39), we get
where the last inequality follows from the condition $\eta\leq\frac{1}{2\sqrt{L_{\mathrm{o}}}}$. This establishes inequality (27) and completes the proof of Lemma 1.
D.1.2 Proof of Lemma 3
Therefore, we have
and find that
Similarly, for the denominator $T_{-}$, we apply equality
It follows that
Then, substituting equations (41a) and (41b) into equation (40), we obtain precisely equation (34) stated in Lemma 3.
D.1.3 Proof of Lemma 4
$\big|\log\mathds{E}[e^{X}]-\mathds{E}[X]\big|=\log\mathds{E}[e^{Y}]\,.$
Finally, applying Hoeffding’s lemma to this resulting log moment-generating function bounds it by at most $2\,\|X\|_{\infty}^{2}$. This establishes the result in Lemma 4.
D.2 Proof of Lemma 2
D.2.1 Overview
$\log\Big(\frac{J_{q}}{1-J_{q}}\Big)-\log\Big(\frac{J_{q}^{\rm old}}{1-J_{q}^{\rm old}}\Big)\;\geq\;\big\{\log\big(\mathds{E}^{+}[\exp(X)]\big)-\log\big(\mathds{E}^{-}[\exp(X)]\big)\big\}-L_{\mathrm{p}}T_{\infty}\cdot\eta^{2}\,$, (42)
where the random variable $X$ is defined as
$\log\big(\mathds{E}^{+}[\exp(X)]\big)-\log\big(\mathds{E}^{-}[\exp(X)]\big)\;\geq\;\Delta\mu_{q}(\uppi_{\theta_{\rm old}})\cdot\eta-\frac{G_{\mathrm{p}}^{2}T_{\psi_{1}}}{1-J_{q}^{\rm old}}\cdot\eta^{2}\,$, (44)
which holds under the condition that
$\eta\;\leq\;\frac{1}{\sqrt{2\,G_{\mathrm{p}}^{2}\cdot T_{\psi_{1}}}}\,.$
Combining the bounds from (42) and (44) directly establishes the result in Lemma 2. We now prove these two intermediate claims in Subsections D.2.2 and D.2.3.
D.2.2 Proof of Inequality (42)
Our starting point is the smoothness property of the log-policy, as stated in Assumption 2. For any prompt-response pair $(q,\vec{\boldsymbol{o}}\!\,)$, the function $\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,\mid q)$ is $(L_{\mathrm{p}}\cdot T_{\infty})$-smooth with respect to $\theta$. A standard result for smooth functions is that:
D.2.3 Proof of Inequality (44)
We now prove each claim in turn.
Proof of Inequality (45):
$\log(u/v)\leq\frac{1}{p}\left(\log(pu+(1-p)v)-\log(v)\right).$
Substituting (48) into (47) completes the proof of inequality (45).
Proof of Inequality (46a):
Substituting this back into our gradient identity yields:
Dividing by $J_{q}^{\rm old}$ and multiplying by $\eta$, we find
Proof of Inequality (46b):
Then $X_{t}=\sum_{t^{\prime}=1}^{t}\xi_{t^{\prime}}$ is a martingale with respect to the filtration $\{\mathcal{F}_{t}\}$, and $X=X_{|\vec{\boldsymbol{o}}\!\,|}$. By Assumption 2, the increments are bounded: $|\xi_{t}|\leq G_{\mathrm{p}}\cdot\eta$.
Supermartingale and Optional Stopping. Using Hoeffding’s lemma (Boucheron et al., 2013, Lemma 2.6) on the conditional expectation, we have
Finally, we bound the remaining term using the sub-exponential ($\psi_{1}$-Orlicz) norm of the response length, $T_{\psi_{1}}$ from Assumption 3. By definition,
$\mathds{E}\big[\exp\{|\vec{\boldsymbol{o}}\!\,|/T_{\psi_{1}}\}\big]\;\leq\;2\,.$
It follows that
Plugging this back into our bound for $\log\mathds{E}[e^{X}]$ gives
$\log\mathds{E}[e^{X}]\;\leq\;G_{\mathrm{p}}^{2}\eta^{2}\cdot T_{\psi_{1}}\,.$
This establishes inequality (46b) and completes the proof.
Appendix E Example of Uniformly Lower Bounded Gap Alignments
To illustrate a situation when the uniformly lower bounded Gap Alignment assumption $\Delta\mu_{q}(k)\geq\Delta\mu_{q}$ of Corollary 3.1 holds, we consider a simplified setting with a softmax policy over two fixed responses: one correct ($\vec{\boldsymbol{o}}\!\,^{+}$) and one incorrect ($\vec{\boldsymbol{o}}\!\,^{-}$).
Setup.
Consider a single prompt $q$ with response space $\mathcal{O}_{q}=\{\vec{\boldsymbol{o}}\!\,^{+},\vec{\boldsymbol{o}}\!\,^{-}\}$, where $r^{\star}(q,\vec{\boldsymbol{o}}\!\,^{+})=1$ and $r^{\star}(q,\vec{\boldsymbol{o}}\!\,^{-})=0$. The policy is parameterized as a softmax over two logits $\theta=(\theta^{+},\theta^{-})\in\mathbb{R}^{2}$:
$\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{+}\mid q)=\frac{e^{\theta^{+}}}{e^{\theta^{+}}+e^{\theta^{-}}}\,,\qquad\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{-}\mid q)=\frac{e^{\theta^{-}}}{e^{\theta^{+}}+e^{\theta^{-}}}\,.$
Thus, the success probability is $J_{q}=\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{+}\mid q)$ and the score functions for each response can be computed as:
$\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{+}\mid q)$ $=\begin{pmatrix}1-J_{q}\\ -(1-J_{q})\end{pmatrix}$
$\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{-}\mid q)$ $=\begin{pmatrix}-J_{q}\\ J_{q}\end{pmatrix}\,$.
In this two-response setting, the conditional distributions $\uppi^{+}_{\theta}(\cdot\mid q)$ and $\uppi^{-}_{\theta}(\cdot\mid q)$ are point masses on $\vec{\boldsymbol{o}}\!\,^{+}$ and $\vec{\boldsymbol{o}}\!\,^{-}$ respectively. Thus: $\boldsymbol{g}_{q}^{+}=\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{+}\mid q)$ and $\boldsymbol{g}_{q}^{-}=\nabla_{\theta}\,\log\uppi_{\theta}(\vec{\boldsymbol{o}}\!\,^{-}\mid q)$. The Gradient Gap is then a constant: $\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}=(1,-1)^{T}$. Thus, for $\boldsymbol{w}_{k}$ chosen in the normalized direction of $\boldsymbol{g}_{q}^{+}-\boldsymbol{g}_{q}^{-}$, we have the Gap Alignment is the constant $\sqrt{2}$.
Appendix F Proofs of Lower Bounds
Intuition for the Lower Bound Analysis.
For the trajectory formulation, our convergence analysis (Theorem 1) relies on equation Equation 14, which uses a first-order approximation of the change in log-odds. For the lower bound, however, it is crucial to examine the second-order expansion. To this end, we define conditional variances over the positive (and negative) response space: ${\rm Var}^{+}:\,={\rm Var}_{\vec{\boldsymbol{o}}\!\,\sim\uppi_{k}(\cdot\mid q,\mathcal{O}_{q}^{+})}\big[\boldsymbol{w}_{k}\cdot\nabla_{\theta}\,\log\uppi_{k}(\vec{\boldsymbol{o}}\!\,\mid q)\big]$. The term $\rm Var$$^{-}$ is defined analogously. The second-order Taylor expansion gives
$\log\Big(\frac{J_{q}(k+1)}{1-J_{q}(k+1)}\Big)-\log\Big(\frac{J_{q}(k)}{1-J_{q}(k)}\Big)\;=\;\Delta\mu_{q}(k)\,\eta_{k}+$ $\{{\rm Var}^{+}-{\rm Var}^{-}\}\cdot\eta_{k}^{2}+\mathcal{O}(\eta_{k}^{3})\,.$
[…]
F.1 Trajectory-Level Analysis Lower Bound (Proof of Theorem 2)